A box of 12 tins of tuna contains 6 which are tainted, suppose 7 tins are opened for inspection, find the probability that (a) exactly 3 of them are tainted, (b) at least 5 of them are tainted, (c) at most 3 of them are tainted


1
Expert's answer
2021-10-28T16:52:04-0400

Let "X=" the number of tainted tins: "X\sim Bin(n, p)."

Given "n=7, p=6\/12=0.5, q=1-p=0.5"

(a)


"P(X=3)=\dbinom{7}{3}(0.5)^3(0.5)^{7-3}=0.2734375"

(b)


"P(X\geq 5)=P(X=5)+P(X=6)+P(X=7)"

"=\dbinom{7}{5}(0.5)^5(0.5)^{7-5}+\dbinom{7}{6}(0.5)^6(0.5)^{7-6}"

"+\dbinom{7}{7}(0.5)^7(0.5)^{7-7}=0.2265625"

(c)


"P(X\leq 3)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)=\dbinom{7}{0}(0.5)^0(0.5)^{7-0}"

"+\dbinom{7}{1}(0.5)^1(0.5)^{7-1}+\dbinom{7}{2}(0.5)^2(0.5)^{7-2}"

"+\dbinom{7}{3}(0.5)^3(0.5)^{7-3}=0.5"